OpenStudy (anonymous):

Solve the compound inequality: -9

OpenStudy (anonymous):

we can have +5 for all :)

OpenStudy (anonymous):

-9+5<x or x<11+5 -4<x or x<16

OpenStudy (anonymous):

separate them, and then solve each of them -9 < x - 5 x - 5 < 11

OpenStudy (anonymous):

Look : If we have : -9+5<x-5+5<11+5 Got it ?

OpenStudy (anonymous):

its -4<x>16

OpenStudy (anonymous):

@lexxilouise : ?????

OpenStudy (anonymous):

So would my solution set be (-4,16)?

OpenStudy (anonymous):

@machaelamae17 : I think when we + numbers to a nagative number will be this u said :)

OpenStudy (anonymous):

-4<x<16

OpenStudy (anonymous):

@machaelamae17 :Thank u but I think your answer is wrong :)

OpenStudy (anonymous):

it is the answer

OpenStudy (anonymous):

what classes do you take @E.ali

OpenStudy (anonymous):

@UditKulka is realy right :) But did u got it @lexxilouise?

OpenStudy (anonymous):

u @machaelamae17 made error i think so

OpenStudy (anonymous):

@UditKulka you think @machaelamae17 is wrong ?

OpenStudy (anonymous):

i just accidently switch the signs its suppose to be -4<x<16

OpenStudy (anonymous):

My online math homework just froze so im not sure yet..

OpenStudy (anonymous):

it was a error

OpenStudy (anonymous):

Is it in interval notation?

OpenStudy (anonymous):

@machaelamae17 Now it s right :) No problem :) Thank u :)

OpenStudy (anonymous):

I think @lexxilouise just need one of our so i go out :)

OpenStudy (anonymous):

Thank you so much everyone!

OpenStudy (anonymous):

i only made a sign mistake next do be as rude and maybe everything will be okay all you had to say was my signs were messed up and i would go back and look at the question there was no reason to be rude and say YOUR WRONG

OpenStudy (anonymous):

\[-18\le5x-14\le1\]

OpenStudy (anonymous):

separate them then solve... -18 <= 5x - 14 -18 + 14 <= 5x -4 <= 5x -4/5 <= x 5x - 14 <= 1 5x <= 1 + 14 5x <= 15 x < = 3

OpenStudy (anonymous):

Is this in interval notation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Would it be [-4/5,3] ??

OpenStudy (anonymous):

yes...use the brackets because there are equal sighs involved.

OpenStudy (anonymous):

Thank you so much! You're the bomb

OpenStudy (anonymous):

lol...I don't know about that. I am glad I coud help you

OpenStudy (anonymous):

@lexxilouise:Do u need help ?:) Got it ?